A 2.30 N object vibrates at the end of a vertical spring. If the frequency of the vibration is 1.00 Hz, what is the spring constant?

This one is giving me trouble.
1
Expert's answer
2020-04-10T08:52:08-0400
"f=\frac{1}{2\pi}\sqrt{\frac{k}{m}}=\frac{1}{2\pi}\sqrt{\frac{gk}{F}}"

"1=\frac{1}{2\pi}\sqrt{\frac{9.8k}{2.3}}"

"k=9.27\frac{N}{m}"